# Diving into a multi-band holographic superconductor

### v2

@@ Abstract
In this paper, we consider the interior structure of a multi-band holographic superconductor model. We focus on the holographic superconductor system with two scalar fields which correspond to two s-wave order parameters in the dual condensed matter system. With two s-wave order parameters, the boundary system has more interesting behaviors which can also be reflected in black hole interior structure. We find the Einstein-Rosen bridge collapse and Josephson oscillations of two scalar fields inside the horizon. For the region near the singularity, we find that the metric still presents Kasner form and there is also Kasner transition behavior. However, when two scalar fields coexist, the Kasner exponents and Kasner transition formula will be different from the single scalar field case. The different interior structures between multi-band holographic superconductor and single-band holographic superconductor we find in this work further confirms that the black hole interior is important to reflect the properties of dual condensed matter systems.
1 Introduction:
@@
AdS/CFT correspondence provides a useful tool to understand the strongly coupled condensed matter systems using dual gravitational framework Maldacena:1997re ; Aharony:1999ti . One prominent example is the holographic superconductor first raised in Ref.Hartnoll:2008vx ; Hartnoll:2008kx which is an s-wave superconductor model. The physical meaning of this holographic superconductor is as follows. From the bulk side, when gauge field and charged scalar fields are added, the original background without scalar hair (Reissner-Nordstrom black hole) will become unstable to form scalar hairs as we tune the temperature or chemical potential of the gravitational system. This spontaneous generated scalar hair can be understood as the order parameter of superconducting phase transition at the boundary by using AdS/CFT dictionary. Based on Ginzburg-Landau theory, this gravitational model mimics the superconducting phase transition. After the first discovery of this s-wave holographic superconductor model, there are many further generalizations to superconductor models with different types of order parameters, such as p-wave superconductor model Gubser:2008wv ; Cai:2013aca and d-wave superconductor model Chen:2010mk ; Benini:2010pr .
@@
However, many realistic high $T_{c}$ superconductor systems are characterized by multi Fermi surfaces which means that there should be multiple order parameters in order to describe these systems PhysRevLett.3.552 ; 10.1143/PTP.29.1 . So many theoretical frameworks have also been established in condensed matter theory to understand multi-band superconductor models which have multiple order parameters PhysRevLett.106.047005 ; PhysRevB.85.134514 . Holography provides a simple way to describe superconductors with multiple order parameters and many holographic models involving coexistence of multiple order parameters have been constructed. Ref.Basu:2010fa builds a simple model to describe multi-band superconductor holographically. In this model, there are two charged scalar fields coupled to the same $U(1)$ gauge field. It has been shown in Ref.Basu:2010fa that there are cases where these two s-wave order parameters coexist. As this work was done in the probe limit, Ref.Cai:2013wma generalized this result to fully back-reacted case and found that the coexistence region will enlarge when increasing the back-reaction. This multi-band model is $s+s$ type, there are also many further generalizations to the coexistence of $s+p$ order parameters Nie:2013sda , $p+p$ order parameters Donos:2013woa and $s+d$ order parameters Li:2014wca . Ref.Cai:2015cya gives a comprehensive review of the development of holographic superconductor model.
@@
Above discussions regarding these holographic superconductor models only consider the black hole exterior region. This is because for the holographic study of condensed matter systems at finite temperature, the boundary state is a mixed state whose gravitational dual is the black hole exterior part. Thus it seems that black hole interior region is not related to these studies. However, the boundary mixed state can be purified to be a pure state. The most well-known purification is canonical purification and the resulting pure state is the thermo-field double state which reads
$|TFD\rangle=\sum_{i}e^{-\beta E_{i}/2}|E_{i}\rangle_{L}|E_{i}\rangle_{R}$ (1)
@@
Once upon purification, the dual gravitational system of this pure state is the eternal AdS black hole (AdS wormhole) which contains the black hole interior part Maldacena:2001kr ; Maldacena:2013xja .For different systems with different Hamiltonians, the energy spectrum $E_{i}$ will be different which leads to different properties of thermo-field double state. Certainly, this will also influence the interior structure of corresponding black holes and thus the features of black hole interior should also be important to illustrate the properties of boundary system.
@@
Based on this consideration, many holographic investigations of boundary system have also been done from black hole interior point of view Cai:2020wrp ; Hartnoll:2020fhc ; Hartnoll:2020rwq ; An:2021plu ; Liu:2021hap ; Cai:2021obq ; An:2022lvo ; Gao:2023zbd ; Cai:2023igv ; Cai:2024ltu ; Carballo:2024hem ; Caceres:2023zhl ; Hartnoll:2022rdv ; Caceres:2021fuw ; Xu:2023fad ; Gao:2023rqc ; Arean:2024pzo ; Caceres:2024edr . For holographic superconductor model, Ref. Hartnoll:2020fhc investigated the interior structure of s-wave holographic superconductor proposed in Hartnoll:2008kx . They found that the black hole interior structure changes drastically after the superconducting phase transition. The original Cauchy horizon disappears due to the presence of scalar hair. And moreover, they also found a series of critical phenomena inside the horizon which are called "ER collapse" and "Josephson oscillation", these phenomena are most prominent near the critical temperature $T_{c}$. For the region near the singularity, there are also new behaviors. The metric takes the form of anisotropic Kasner universe and for unstable Kasner region, there also exists Kasner inversion behavior which follows the BKL transition rule Belinsky:1970ew . It has also been found that different boundary models have different interior structures, for example the s-wave holographic superconductor model with exponential-type scalar potential and p-wave holographic superconductor model all bear distinct interior structures Cai:2021obq ; An:2022lvo ; Cai:2023igv ; Cai:2024ltu . Thus the black hole interior structures are very useful to characterize the distinct features of the boundary model.
@@
Besides these progress, the investigations regarding the interior structure all focus on the single order parameter case. There is still no research on the interior structure of multi-band holographic superconductor model. From condensed matter point of view, the physics of multi-band superconductor will be very different from single-band superconductor, understanding this difference holographically is a very interesting question. As the black hole interior structure can be a sensitive probe to the boundary model, thus we will investigate the interior structure of multi-band holographic superconductor in this paper. For simplicity, we will firstly focus on the multi-band holographic superconductor with two charged scalar fields. We find that with one additional scalar field, the Josephson oscillation behavior, Kasner geometry and Kasner transition behavior will all bear new features compared to single scalar field case.
@@
The structure of this paper is as follows. In Sec.2, we first review the multi-band holographic superconductor with two competing scalar fields which have different mass and charge. We will show the three different types of coexistence behavior for different choice of parameters. In Sec.3, we will choose one specific coexistence scenario as a representative and investigate the interior structure of this hairy black hole with multiple scalar fields. We show the ER collapse, Josephson oscillation and Kasner behavior inside this black hole and describe its difference with the single scalar field case. In Sec.4, we conclude our paper and give some future directions.
2 Multi-band holographic superconductor with two scalar fields:
@@
In this section, we will first give a brief introduction to the holographic multi-band superconductor model where two order parameters can coexist and compete. We will only focus on the case where two competing condensates are both scalar. To model the two s-wave orders on the boundary, we need to introduce two scalar fields in the bulk. Thus the action takes following form
$S=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}[R+\frac{6}{L^{2}}-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\sum_{k=1}^{2}(-|\nabla\psi_{k}-ie_{k}A\psi_{k}|^{2}-m_{k}^{2}|\psi_{k}|^{2})]$ (2)
@@
where the two scalar fields are minimally coupled to the same $U(1)$ gauge field and there is no direct interaction between them. $e_{k}$ and $m_{k}$ are the charge and mass of scalar field $\psi_{k}$, $\kappa^{2}=8\pi G$ is the Planck length and $F_{\mu\nu}=\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}$ is the electro-magnetic field strength.
Varying the action with respect to the metric field $g_{\mu\nu}$, electromagnetic field $A_{\mu}$ and scalar field $\psi_{k}$, we found the equation of motion reads,
$D_{1\mu}D_{1}^{\mu}\psi_{1}-m_{1}^{2}\psi_{1}=0$ (3)
$D_{2\mu}D_{2}^{\mu}\psi_{2}-m_{2}^{2}\psi_{2}=0$ (4)
$\nabla^{\mu}F_{\mu\nu}=ie_{1}[\psi_{1}^{*}D_{1\mu}\psi_{1}-\psi_{1}D_{1\mu}^{*}\psi_{1}^{*}]+ie_{2}[\psi_{2}^{*}D_{2\mu}\psi_{2}-\psi_{2}D_{2\mu}^{*}\psi_{2}^{*}]$ (5)
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}-\frac{3}{L^{2}}g_{\mu\nu}=\frac{1}{2}F_{\mu\lambda}F_{\nu}^{\lambda}+\frac{1}{2}[(D_{1\mu}\psi_{1}D_{1\nu}^{*}\psi_{1}^{*}+D_{2\mu}\psi_{2}D_{2\nu}^{*}\psi_{2}^{*})+\mu\leftrightarrow\nu]+\frac{g_{\mu\nu}}{2}L_{m}$ (6)
where $L_{m}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-|D_{1}\psi_{1}|^{2}-m_{1}^{2}|\psi_{1}|^{2}-|D_{2}\psi_{2}|^{2}-m_{2}^{2}|\psi_{2}|^{2}$ and we have defined $D_{1\mu}=\nabla_{\mu}-ie_{1}A_{\mu}$, $D_{2\mu}=\nabla_{\mu}-ie_{2}A_{\mu}$.
We consider the static and spherically symmetric metric ansatz
$ds^{2}=\frac{1}{z^{2}}(-fe^{-\chi}dt^{2}+\frac{dz^{2}}{f}+dx^{2}+dy^{2})$ (7)
and corresponding matter field ansatz
@@
With these special ansatz, the equations of motion simplifies to
$z^{2}e^{\chi/2}(zh\psi_{1}^{\prime})^{\prime}=\frac{m_{1}^{2}\psi_{1}}{z^{2}}-\frac{e^{\chi/2}e_{1}^{2}\phi^{2}\psi_{1}}{hz^{3}}$ (9)
$z^{2}e^{\chi/2}(zh\psi_{2}^{\prime})^{\prime}=\frac{m_{2}^{2}\psi_{2}}{z^{2}}-\frac{e^{\chi/2}e_{2}^{2}\phi^{2}\psi_{2}}{hz^{3}}$ (10)
$z^{2}(e^{\chi/2}\phi^{\prime})^{\prime}=\frac{2\phi(e_{1}^{2}\psi_{1}^{2}+e_{2}^{2}\psi_{2}^{2})}{hz^{3}}$ (11)
$\chi^{\prime}=z(\psi_{1}^{\prime 2}+\psi_{2}^{\prime 2})+\frac{\phi^{2}}{h^{2}z^{5}}(e_{1}^{2}\psi_{1}^{2}+e_{2}^{2}\psi_{2}^{2})$ (12)
$4z^{4}e^{\chi/2}h^{\prime}=-12+2m_{1}^{2}\psi_{1}^{2}+2m_{2}^{2}\psi_{2}^{2}+e^{\chi}z^{4}\phi^{\prime 2}$ (13)
where $h=\frac{fe^{-\chi/2}}{z^{3}}$ is introduced for convenience.The first two equations are scalar field equations and the third equation is Maxwell equation. The last two equations are two independent Einstein equations. Solving these five coupled differential equations with suitable boundary condition will holographically model the multi-band superconductor system. The parameter $e_{1}$,$e_{2}$ controls the back-reaction of the scalar field to the background. It has been found that the strength of backreaction will strongly influence the coexistence behavior of two scalar order parameter. To be more specific, increasing the backreaction ( decreasing the charge) will enlarge the coexistence regionCai:2013wma .  [^2] Note that infinite charge limit corresponds to probe limit where back-reaction effect is switched off
For the convenience to perform numerics, we choose mass parameter of two scalar fields to be $m_{1}^{2}=0$ and $m_{2}^{2}=-2$ respectively. For the spacetime having asymptotic AdS boundary, the general falloff of matter fields and metric fields near the boundary takes the form
$\psi_{1}=\psi_{1}^{(0)}z^{\Delta_{1-}}+\psi_{1}^{(1)}z^{\Delta_{1+}}+...,\quad\psi_{2}=\psi_{2}^{(0)}z^{\Delta_{2-}}+\psi_{2}^{(1)}z^{\Delta_{2+}}+...$ (14)
@@
$\psi_{1}=\psi_{1}^{(0)}+\psi_{1}^{(1)}z^{3}+...,$ (16)
$\psi_{2}=\psi_{2}^{(0)}z+\psi_{2}^{(1)}z^{2}+...$ (17)
We choose standard quantization procedure in this paper where $\psi^{(0)}$ is identified as the source and $\psi^{(1)}$ as the expectation value. To build a superconductor model, we require the scalar condensate is spontaneously generated thus we should set the boundary source term to be zero $\psi_{1,2}^{(0)}=0$. By expanding the field near the horizon and plugging in the equations of motion, one finds that there are five independent parameters at the horizon $\{z_{h},\psi_{1}(z_{h}),\psi_{2}(z_{h}),\phi^{\prime}(z_{h}),\chi(z_{h})\}$. Note that there are scaling symmetries associated to the equations of motion
@@
When performing numerics, we can first set $z_{h}=1$ and $\chi(z_{h})=0$ and then use the first scaling symmetry to set $\chi(0)=0$, this condition guarantees the asymptotic condition where the boundary time is the same as bulk time. Then we only have three independent parameters $\{\psi_{1}(z_{h}),\psi_{2}(z_{h}),\phi^{\prime}(z_{h})\}$, two of them will be chosen as shooting parameters to match the condition $\psi_{1,2}^{(0)}=0$. Moreover for the expansion of electromagnetic field at boundary in Eq.(14), $\mu$ is the chemical potential and $\rho$ is the charge density. We will work in grand canonical ensemble where the chemical potential should be kept fixed. By using the second scaling symmetry , we can set $\mu=1$ at the boundary to achieve this.
@@
At high temperature, the normal phase is dominant which is RN-AdS black hole where two scalar field vanishes $\psi_{1}=\psi_{2}=0$. After decreasing temperature, the spontaneous symmetry breaking happens and scalar order parameter appears. It is interesting to note that when there are two scalar fields, when further lowering the temperature, second order parameter will appear which have significant influence on the first order parameter. Most notably, it was plotted in Fig.1 that second order parameter will suppress the first order parameter  [^3] Note that we rescale the condensate value using $e_{2}$ in this figure to match the notation in Ref.Cai:2013wma . and by further decreasing temperature, the second order parameter will still increase and the first order parameter will decrease to zero. This case is called phase-X in Ref.Cai:2013wma .
@@
Figure 1: Phase X of multiple order parameters in holographic superconductor. This figure is plotted for parameter $e_{2}=4$ and $e_{1}/e_{2}=1.95$.The red curve corresponds to $\psi_{1}$ while the blue curve corresponds to $\psi_{2}$. We fix chemical potential $\mu=1$ in this figure.It can be seen that when the second order parameter appears, it suppresses the first order parameter
In this case, the two order parameters have strong competition between each other. The competition can also be seen in black hole exterior configuration shown in Fig.2.
Figure 2: The configuration of charge black holes with scalar hair from the event horizon, $z=z_{H}$, to the AdS boundary, $z=0$, at different temperatures. We have fixed the chemical potential $\mu=1$.
@@
Furthermore, by increasing the backreaction (decreasing the parameter $e_{1}$ or $e_{2}$), it was found in Ref.Cai:2013wma that strong backreaction will enlarge the coexistence region which makes two order parameters always coexist. Ref.Cai:2013wma found two possible phases, one is shown in the left panel of Fig.3 for the parameter $e_{2}=2$ and $e_{1}/e_{2}=1.95$ which is called phase-A in Cai:2013wma , and another is shown in the right panel of Fig.3 for the parameter choice $e_{2}=1.5$ and $e_{1}/e_{2}=1.9$ which is called phase-B in Cai:2013wma .
@@
Figure 3: Multiple order parameters in holographic superconductor for larger backreaction. Left panel: Phase A with $e_{2}=2$ and $e_{1}/e_{2}=1.95$. Right panel: Phase B with $e_{2}=1.5$ and $e_{1}/e_{2}=1.9$. The red curve corresponds to $\psi_{1}$ while the blue curve corresponds to $\psi_{2}$. We fix chemical potential $\mu=1$ in this figure. It can be seen that by increasing the backreaction, the suppressive effect of second order parameter will be less dramatic and the two order will always coexist.
@@
These condensate behaviors imply that the microscopic physics behind this superconductor model will be much more complex than the single order parameter case. It is natural to expect that this complicated boundary physical structure can also be reflected in terms of black hole interior. Thus in the next section, we will investigate this physical behavior from the black hole interior point of view.
3 Interior structure of multi-band holographic superconductor:
In this section, we will investigate the interior structure of multi-band holographic superconductor We only investigate these phenomena in Phase-X case, behaviors in Phase-A and Phase-B case are similar.

### v3

# Diving into a holographic multi-band superconductor

@@ Abstract
In this work, we investigate the interior structure of a holographic multi-band superconductor model. We focus on the holographic superconductor system with two scalar fields which correspond to two s-wave order parameters in the dual condensed matter system. We discover a new kind of transition rule among Kasner universe near the black hole singularity which is distinct from the holographic single-band superconductor model. This transition rule is the first generalization of Kasner transition behavior to scenarios involving multiple free parameters, which is useful for uncovering the most general interior structures of hairy black holes. Moreover, we find that the Kasner exponents are sensitive to the details of order parameters in boundary system. These different near singularity structures we found show that the black hole interior plays crucial role in characterizing the boundary dual condensed matter systems.
keywords:
holographic duality , black hole , singularity
journal: Physics Letters B
[…]
1 Introduction
@@
AdS/CFT correspondence provides a useful tool to understand the strongly coupled condensed matter systems using dual gravitational framework [1, 2]. One prominent example is the holographic superconductor first raised in Ref.[3, 4] which is an s-wave superconductor model. The physical meaning of this holographic superconductor is as follows. From the bulk side, when gauge field and charged scalar fields are added, the original background without scalar hair (Reissner-Nordstrom black hole) will become unstable to form scalar hairs as we tune the temperature or chemical potential of the gravitational system. This spontaneous generated scalar hair can be understood as the order parameter of superconducting phase transition at the boundary by using AdS/CFT dictionary. Based on Ginzburg-Landau theory, this gravitational model mimics the superconducting phase transition. After the first discovery of this holographic s-wave superconductor model, there are many further generalizations to superconductor models with different types of order parameters, such as p-wave superconductor model [5, 6] and d-wave superconductor model [7, 8].
@@
However, instead of one order parameter, investigations on many realistic superconductor systems (such as $MgB_{2}$ and $Sr_{2}RuO_{4}$ etc) should include multiple order parameters [9, 10]. Many theoretical frameworks have also been established in condensed matter theory to understand multi-band superconductor models which have multiple order parameters [11, 12]. Besides these theories, AdS/CFT correspondence can also provide an elegant method to describe superconductors with the coexistence of multiple order parameters. Ref.[13] builds a simple model to describe multi-band superconductor holographically. In this model, there are two charged scalar fields which are coupled to the same $U(1)$ gauge field. It has been shown in Ref.[13] that there are cases where these two s-wave order parameters coexist. As this work was done in the probe limit, Ref.[14] generalized this result to fully back-reacted case and found that the coexistence region will enlarge when increasing the back-reaction. This multi-band model is $\mathrm{s+s}$ type, there are also further generalizations to the coexistence of $\mathrm{s+p}$ order parameters [15, 16], $\mathrm{p+p}$ order parameters [17] and $\mathrm{s+d}$ order parameters [18]. Ref.[19] gives a comprehensive review of these holographic multi-band superconductor models.
@@
Above references regarding these holographic superconductor models only focus on the black hole exterior region. This is because for the holographic study of condensed matter systems at finite temperature, the boundary state is a mixed state whose gravitational dual is the black hole exterior part. Thus it seems that black hole interior region is not related to these studies. However, the boundary mixed state can be purified to be a pure state. The most well-known purification is canonical purification and the resulting highly entangled pure state is the thermo-field double state which reads
$|TFD\rangle=\frac{1}{\sqrt{Z}}\sum_{i}e^{-\beta E_{i}/2}|E_{i}\rangle_{L}|E_{i}\rangle_{R}.$ (1)
@@
where $Z=\sum_{i}e^{-\beta E_{i}}$. Once upon purification, the dual gravitational system of this pure state is the eternal AdS black hole (AdS wormhole) which contains the black hole interior part [20, 21]. For different systems with different Hamiltonians, the energy spectrum $E_{i}$ will be different which leads to different thermo-field double state. Certainly, this will also influence the interior structure of corresponding black holes and thus the features of black hole interior should also be important to illustrate the properties of boundary system.
@@
Based on this consideration, many holographic investigations of boundary system have also been done from black hole interior point of view [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. For holographic superconductor model, Ref. [23] investigated the interior structure of s-wave holographic superconductor proposed in [4]. They found that the black hole interior structure changes drastically after the superconducting phase transition. The original Cauchy horizon disappears due to the presence of scalar hair. And moreover, they also found a series of critical phenomena inside the horizon which are called ”ER collapse” and ”Josephson oscillation”, these phenomena are most prominent near the critical temperature $T_{c}$. For the region near the singularity, there are also new behaviors. The metric takes the form of anisotropic Kasner universe and there can also exist transitions among these Kasner universe which is similar to the well-known BKL transition [40, 41, 42]. More notably, black hole interior structures are more sensitive to the features of boundary theories which is different from the the black hole exterior. For example, the s-wave holographic superconductor model with exponential-type scalar potential and p-wave holographic superconductor model all bear distinct interior structures [27, 28, 30, 31]. Therefore, the black hole interior structures are very important in the holographic descriptions of condensed matter systems.
@@
For the above progress, the investigations regarding the interior structure all focus on the single order parameter case. There is still no research on the interior structure of holographic multi-band superconductor model. From condensed matter physics, the properties of multi-band superconductor will be very different from single-band superconductor, thus how to characterize this difference holographically is a very interesting question. As the black hole interior structure is a sensitive probe to the boundary model which depends on the models significantly, investigating interior structures of the holographic multi-band superconductor is well motivated which is the main topic of this paper. For simplicity, we will firstly focus on the holographic multi-band superconductor with two charged scalar fields. We find that by incorporating multiple scalar fields, the Kasner geometry and Kasner transition behavior near the black hole singularity will present new features compared to single scalar field case. These new features are not only important to characterize the properties of multi-band superconductor but also constitute an important part towards finding the most general interior structure of hairy black holes.
@@
The structures of this letter are as follows. In Sec.2, we first review the holographic multi-band superconductor with two competing scalar fields which have different mass and charge. We will show the three different types of coexistence behavior for different choice of charge parameters. In Sec.3, we will investigate the interior structure of this holographic multi-band superconductor. We investigate both the near horizon region and near singularity region and describe its difference between the holographic single-band superconductor model. In Sec.4, we conclude our paper and give some future directions.
2 Holographic multi-band superconductor with two scalar fields:
@@
In this section, we will first give a brief introduction to the holographic multi-band superconductor model where two order parameters can coexist and compete. We will only focus on the case where two competing condensates are both scalar. To model the two s-wave orders on the boundary, two scalar fields are introduced in the bulk theory. Thus the action takes following form
$S=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}[R+\frac{6}{L^{2}}-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\sum_{k=1}^{2}(-|\nabla\psi_{k}-ie_{k}A\psi_{k}|^{2}-m_{k}^{2}|\psi_{k}|^{2})],$ (2)
@@
where the two scalar fields are minimally coupled to the same $U(1)$ gauge field and there is no direct coupling between them. $e_{k}$ and $m_{k}$ are the charge and mass of scalar field $\psi_{k}$, $\kappa^{2}=8\pi G$ is the Planck length and $F_{\mu\nu}=\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}$ is the electro-magnetic field strength.
For time-independent s-wave superconductor model, the static and spherically symmetric metric ansatz should be imposed as
$ds^{2}=\frac{1}{z^{2}}(-fe^{-\chi}dt^{2}+\frac{dz^{2}}{f}+dx^{2}+dy^{2}),$ (3)
moreover, the corresponding matter field ansatz read
@@
With these special ansatz, by varying the action with respect to the scalar field $\psi_{k}$, electromagnetic field $A_{\mu}$ and the metric field $g_{\mu\nu}$, the equations of motion read
$z^{2}e^{\chi/2}(zh\psi_{1}^{\prime})^{\prime}=\frac{m_{1}^{2}\psi_{1}}{z^{2}}-\frac{e^{\chi/2}e_{1}^{2}\phi^{2}\psi_{1}}{hz^{3}}\ ,$ (5)
$z^{2}e^{\chi/2}(zh\psi_{2}^{\prime})^{\prime}=\frac{m_{2}^{2}\psi_{2}}{z^{2}}-\frac{e^{\chi/2}e_{2}^{2}\phi^{2}\psi_{2}}{hz^{3}}\ ,$ (6)
$z^{2}(e^{\chi/2}\phi^{\prime})^{\prime}=\frac{2\phi(e_{1}^{2}\psi_{1}^{2}+e_{2}^{2}\psi_{2}^{2})}{hz^{3}}\ ,$ (7)
$\chi^{\prime}=z(\psi_{1}^{\prime 2}+\psi_{2}^{\prime 2})+\frac{\phi^{2}}{h^{2}z^{5}}(e_{1}^{2}\psi_{1}^{2}+e_{2}^{2}\psi_{2}^{2})\ ,$ (8)
$4z^{4}e^{\chi/2}h^{\prime}=-12+2m_{1}^{2}\psi_{1}^{2}+2m_{2}^{2}\psi_{2}^{2}+e^{\chi}z^{4}\phi^{\prime 2}\ ,$ (9)
where $h=\frac{fe^{-\chi/2}}{z^{3}}$ is introduced for convenience. The first two equations are scalar field equations and the third equation is Maxwell equation. The last two equations are two independent Einstein equations. Solving these five coupled differential equations with suitable boundary condition will holographically model the multi-band superconductor system. The parameter $e_{1}$,$e_{2}$ controls the back-reaction of the scalar field to the background. It has been found that the strength of back-reaction will strongly influence the coexistence behavior of two scalar order parameters. To be more specific, increasing the back-reaction (decreasing the charge) will enlarge the coexistence region [14].  [^2] Note that infinite charge limit corresponds to probe limit where back-reaction effect is switched off.
For the convenience to perform numerics, mass parameters of two scalar fields are chosen to be $m_{1}^{2}=0$ and $m_{2}^{2}=-2$. For the spacetime having asymptotic AdS boundary, the general falloff of matter fields and metric fields near the boundary takes the form
$\psi_{1}=\psi_{1}^{(0)}z^{\Delta_{1-}}+\psi_{1}^{(1)}z^{\Delta_{1+}}+...\ ,\quad\psi_{2}=\psi_{2}^{(0)}z^{\Delta_{2-}}+\psi_{2}^{(1)}z^{\Delta_{2+}}+...\ ,$ (10)
@@
$\psi_{1}=\psi_{1}^{(0)}+\psi_{1}^{(1)}z^{3}+...\ ,\quad\psi_{2}=\psi_{2}^{(0)}z+\psi_{2}^{(1)}z^{2}+...\ .$ (12)
Standard quantization procedure is chosen in this paper where $\psi^{(0)}$ is identified as the source and $\psi^{(1)}$ as the expectation value. To build a superconductor model, scalar condensate should be spontaneously generated thus the boundary source terms are set to be zero
$\psi_{1}^{(0)}=0,\quad\psi_{2}^{(0)}=0.$ (13)
To avoid coordinate singularity at horizon, one can expand the fields near the horizon and plug the series into the equations of motion. It can be found that there are five independent parameters at the horizon $\{z_{h},\psi_{1}(z_{h}),\psi_{2}(z_{h}),\phi^{\prime}(z_{h}),\chi(z_{h})\}$. Note that there are scaling symmetries associated to the equations of motion
@@
When performing numerics, one can first set $z_{h}=1$ and $\chi(z_{h})=0$ and then use the first scaling symmetry to set $\chi(0)=0$,  [^3] This condition guarantees the asymptotic condition where the boundary time is the same as bulk time. thus only three independent parameters $\{\psi_{1}(z_{h}),\psi_{2}(z_{h}),\phi^{\prime}(z_{h})\}$ remain. For this boundary value problem, we should invoke the shooting method to numerically solve the differential equations (5-9) with boundary conditions (13). The specific procedure of numerical computation is as follows: firstly under prescribed $\phi^{\prime}(z_{h})$, the fourth-order Runge-Kutta method was selected to numerically determine the boundary value of $\psi_{1}^{(0)}$ and $\psi_{2}^{0}$ under specified choice of shooting parameters $\{\psi_{1}(z_{h}),\psi_{2}(z_{h})\}$. If the boundary conditions are not satisfied, we use the Newton-Raphson iterative method to search for $\{\psi_{1}(z_{h}),\psi_{2}(z_{h})\}$ which matches boundary conditions. After getting the parameters on the horizon, the coupled differential equations (5-9) are directly solved and the expectation value of order parameter $\psi_{1,2}^{(1)}$ are extracted.
@@
Moreover, for the expansion of electromagnetic field at boundary in Eq.(10), $\mu$ is the chemical potential and $\rho$ is the charge density. We choose to work in grand canonical ensemble thus the chemical potential should be kept fixed. We will fix the chemical potential $\mu=1$ by using the second scaling symmetry Eq.(15) in this work.
@@
From the above prescriptions, the condensate behaviors can be deduced. At high temperature, the normal phase is dominant which is RN-AdS black hole where two scalar field vanishes $\psi_{1}=\psi_{2}=0$. After decreasing temperature, the spontaneous symmetry breaking happens and scalar order parameter appears. It is interesting to note that when there are two scalar fields, when further lowering the temperature, second order parameter will appear which have significant influence on the first order parameter. Most notably, it was plotted in Fig.1 that second order parameter will suppress the first order parameter  [^4] Note that we rescale the condensate value using $e_{2}$ in this figure to match the notation in Ref.[14]. and by further decreasing temperature, the second order parameter will still increase and the first order parameter will decrease to zero. This case is called phase-C in Ref.[14].
@@
Figure 1: Phase C of multiple order parameters in holographic superconductor. This figure is plotted for parameter $e_{2}=4$ and $e_{1}/e_{2}=1.95$.The red curve corresponds to $\psi_{1}$ while the blue curve corresponds to $\psi_{2}$. We fix chemical potential $\mu=1$ in this figure.It can be seen that when the second order parameter appears, it suppresses the first order parameter.
@@
In this case, the two order parameters have strong competition between each other. Furthermore, by increasing the back-reaction (decreasing the parameter $e_{1}$ or $e_{2}$), it was found in Ref.[14] that strong back-reaction will enlarge the coexistence region which makes two order parameters always coexist. Ref.[14] found two possible phases, one is shown in the left panel of Fig.2 for the parameter $e_{2}=2$ and $e_{1}/e_{2}=1.95$ which is called phase-A in [14] , and another is shown in the right panel of Fig.2 for the parameter choice $e_{2}=1.5$ and $e_{1}/e_{2}=1.9$ which is called phase-B in [14].
@@
Figure 2: Multiple order parameters in holographic superconductor for larger back-reaction. Left Panel:Phase A with $e_{2}=2$ and $e_{1}/e_{2}=1.95$. Right Panel:Phase B with $e_{2}=1.5$ and $e_{1}/e_{2}=1.9$. The red curve corresponds to $\psi_{1}$ while the blue curve corresponds to $\psi_{2}$. We fix chemical potential $\mu=1$ in this figure. It can be seen that by increasing the back-reaction, the suppressive effect of second order parameter will be less dramatic and the two order will always coexist.
@@
These condensate behaviors imply that the microscopic physics behind this superconductor model will be much more complex than the single order parameter case. Due to the close connection between black hole interior structure and boundary system, it is natural to expect that the coexistence of two order parameters can also significantly influence the black hole interior. Thus in the next section, we will investigate the corresponding black hole interior structure to show this distinctiveness.
3 Interior structure of holographic multi-band superconductor:
By directly solving the equations of motion with specific shooting parameters, the condensate behaviors and exterior region of holographic superconductor can be known. However, besides the exterior part, it is also interesting to directly go beyond the horizon to solve the interior region of holographic superconductor which presents many interesting features [24]. Therefore, we will investigate the interior structure of holographic multi-band superconductor in this section. For the purpose of simplicity, we only investigate these phenomena in Phase-C case, behaviors in Phase-A and Phase-B case are similar.